### v1

Abstract
Abstract: Vineyards are a common way to study persistence diagrams of a data set which is changing, as strong stability means that it is possible to pair points in “nearby” persistence diagrams, yielding a family of point sets which connect into curves when stacked. Recent work has also studied monodromy in the persistent homology transform, demonstrating some interesting connections between an input shape and monodromy in the persistent homology transform for 0-dimensional homology embedded in $\mathbb{R}^{2}$. In this work, we re-characterize monodromy in terms of periodicity of the associated vineyard of persistence diagrams.
We construct a family of objects in any dimension which have non-trivial monodromy for $l$-persistence of any periodicity and for any $l$. More generally we prove that any knot or link can appear as a vineyard for a shape in $\mathbb{R}^{d}$, with $d\geq 3$. This shows an intriguing and, to the best of our knowledge, previously unknown connection between knots and persistence vineyards. In particular this shows that vineyards are topologically as rich as one could possibly hope.
keywords:
Monodromy, Knots and links, Persistent homology, Vineyards
1 Introduction
Computational topology has a number of different branches which, despite some overlap, have remained fairly distinct in nature. On the one hand there is persistent homology [18, 22, 16], a more recent and active area which has its roots in algebraic topology; see [21, 14] for recent survey books on this active topic. On the other hand there are the computational aspects of knot theory and the study and characterization of low dimensional manifolds’ topology, which has a long history of algorithmic development, perhaps dating back to Dehn’s algorithm [12, 13] and the many following algorithmic results in more recent decades on shape and knot recognition in low dimensions [23, 27, 19, 5, 6].
The first goal of this paper is studying an interesting new link between the these two important branches of computational topology. Thanks to the work of Alexander [2] we know that every knot or link can be represented as a braid, that is, for every link there is a braid such that if we glue the ends of the braid, we recover the link. Recall also that for a continuous one parameter family of filtrations, we can “stack” the persistence diagrams of these filtrations; we call the resulting object a vineyard [11]. Thanks to the stability of persistence diagrams, the points in the persistence diagram move continuously (even Lipschitz continuously) with the parameter. This means that we can follow a point in (the stack of) the persistence diagrams; the resulting curve is called a vine. We will prove that for every link there exists an embedded manifold and a family of functions on $\mathcal{M}$ (where each function is induced by the distance to a point in the ambient space, and where in turn each point comes from a curve $\gamma$) such that the vineyard of the family of function yields the braid representing the knot or link in the sense of Alexander.
The second goal of this paper is to show that any type of monodromy can occur in vineyards. This is part of a new research direction in computational topology. In [3], the occurrence of monodromy in the context of the directional persistence transform is studied, more precisely for 0-dimensional persistence modules of objects embedded in $\mathbb{\mathbb{R}}^{2}$. They conclude with an open question about demonstrating monodromy in higher dimensions, as well as several interesting and more open ended questions related to better understanding what monodromy is capturing about the input shape.
1.1 Our contributions
In this paper we exhibit that any type of monodromy and the braid associated to any knot or link can occur in a vineyard. To make our statement more precise, however, we need to introduce some nomenclature, although full definitions will be deferred until Section 2.
Monodromy is the effect where if one makes a loop in a base space of a covering or fibre bundle, the lifted curve may not end up in the same point as you started out with. We say the monodromy is of period $2\pi k$ (with $k>0$) if the lifted curve returns to starting point after $k$ revolutions in the base space. In our context the base space is a closed curve or loop $\gamma:[0,2\pi]\to\mathbb{R}^{d}$ into which we have embedded a manifold $\mathcal{M}$ (mostly some knot, link or some offset of a knot or link). The fibres are the persistence diagrams of the distance function restricted to the manifold $\mathcal{M}$, that is $d(x,\gamma(t))_{\mathcal{M}}$. The bundle therefore is the vineyard. The lifted curve is a vine $\tilde{\gamma}_{(b_{0},d_{0})}(t)$ in the vineyard starting at $(b_{0},d_{0})$ in the persistence diagram of $d(x,\gamma(0))_{\mathcal{M}}$ and the periodicity is the smallest $k>0$, such that for all $i$, $\tilde{\gamma}_{(b_{0},d_{0})}(0)=\tilde{\gamma}_{(b_{i},d_{i})}(2\pi k)$, where we assume that the vine is non-degenerate in the sense that it stays away from the diagonal.
More precisely the two main statements are:
Theorem 1.1.
The persistence distance transform in $\mathbb{R}^{d}$ can exhibit monodromy for persistence up to the ($d-2$)th homology and for extended persistence up to the ($d-1$)th homology. Moreover the periodicity of the monodromy can be $2k\pi$ for any $k\in\mathbb{Z}_{\geq 2}$.
Figure 1: For a given a link (top), consisting of a trefoil and a circle, we construct an angular braid embedded near an annulus (bottom left), which we then modify (bottom right) such that vineyard consisting of the persistence diagrams of the distance function to a point following the twisted annulus on the ‘outside’ contains the input braid with some surgery. See also Figure 2 and the corresponding Example in Appendix A.
Figure 2: Top: side view of the embedding of the link depicted in Figure 1 following a twisted annulus, see also the proof of Theorem 1.1 and the corresponding Example in Appendix A. The observation loop is shown in light blue, with the front part darkened to correspond with the period shown in the bottom figure. Middle: Front-angled view of the embedded link. Bottom: the fraction of the full period (from 00 to $2\pi$) depicted in dark blue above of the braided vineyard of this link, which captures all crossing in the vineyard and exhibits monodromy of period $2\pi\cdot 3$. To transform this braided vineyard into a closed braid, identify the “sides”, or slices, at 00 and $2\pi$.
Theorem 1.2.
Given a knot or link, $d,l\in\mathbb{Z}_{>0}$, with $d\geq 3$ and $l<d-2$, then there exists an $\mathcal{M}\subset\mathbb{R}^{d}$ and a closed curve $\gamma\subset\mathbb{R}^{d}$ such that by identifying the ends of the $l$-vineyard of $d(x,\gamma(t))_{\mathcal{M}}$ will yield a knot or link, which contains the given knot or link as a subset, that is, it is topologically the knot or link we started our construction with after removing some spurious connected components.
It is worth emphasizing that in our constructive proof, the choice of not only $\mathcal{M}$ but also $\gamma$ is critical, as there are choices of $\gamma$ which yield no links in the vineyard.
2 Preliminaries
2.1 Monodromy
Monodromy is an important concept in mathematics that appears in various guises. We refer to the review [15] (the first part of which is almost a review of reviews) and the other reviews mentioned in that paper for an overview of the various aspects of the theory. In this paper we will only consider the simplest incarnation, and only in the setting of topology data analysis.
Let $\tilde{X}$ be a covering space of $X$ with covering map $C:\tilde{X}\to X$, that is for every $x\in X$ there exists an open neighbourhood $x\in U$ and a discrete set $J$, such that $C^{-1}(U)=\sqcup_{i\in J}V_{i}$ and $C|_{V_{j}}:V_{j}\to U$ is a homeomorphism for all $j\in J$. We call the inverse images of points $x\in X$ of the map $C$ the fibres. For a curve $\gamma:[0,2\pi]\to X$ we write $\tilde{\gamma}$ for (one of) its lift(s), that is a continuous map $\tilde{\gamma}:[0,2\pi]\to\tilde{X}$ such that $C\cdot\tilde{\gamma}=\gamma$. If $\gamma$ is a loop, that is $\gamma(0)=\gamma(2\pi)$, then we say that $\gamma$ exhibits monodromy (at the starting point $\tilde{\gamma}(0)\in C^{-1}(\gamma(0))$) if we have that the start and end points of its lift $\tilde{\gamma}$ are different, i.e. $\tilde{\gamma}(0)\neq\tilde{\gamma}(2\pi)$. The difference between $\tilde{\gamma}(0)$ and $\tilde{\gamma}(2\pi)$ is also referred to as monodromy (this difference can in certain cases be best represented by a group, see [15], although we will not need this in our discussion).
Figure 3: Here we see a cover $\tilde{X}$ (in this case a double cover) of the base space $X$, in this case a circle as well as the curve $\gamma$ and its lift $\tilde{\gamma}$.
Because $\gamma$ is a loop we can extend it, formally speaking by concatenating with itself. Here we adopt the convention that the if $\gamma$ and $\tilde{\gamma}$ are two curves parametrized by $[0,2\pi]$, then the concatenation $\gamma\circ\tilde{\gamma}$ is parametrized by $[0,4\pi]$, that is we don’t rescale the parametrization interval. We write
$\gamma^{k}=\underbrace{\gamma\circ\dots\circ\gamma}_{k}$
and $\widetilde{\gamma^{k}}$ for its lifting. We stress that generally
$\widetilde{\gamma^{k}}\neq\underbrace{\tilde{\gamma}\circ\dots\circ\tilde{\gamma}}_{k},$
where $\tilde{\gamma}$ is the lifting of $\gamma$. In fact the right hand side does not even have to be a continuous curve. We say that a loop $\gamma$ (parametrized by $[0,2\pi]$) in the base space $X$ exhibits monodromy of order $k$ if $k$ is the smallest positive integer such that the lifted curve $\widetilde{\gamma^{k}}$ satisfies
$\widetilde{\gamma^{k}}(0)=\widetilde{\gamma^{k}}(2\pi k).$
If $k=1$ we say that $\gamma$ exhibits no or trivial monodromy.
2.2 Knots, links, and braids
In this subsection we briefly recall the formal definitions of knots, links and braids. An (oriented) knot is the the equivalence class of oriented closed curves embedded in $3$-dimensional Euclidean space, $\gamma\colon\mathbb{S}^{1}\to\mathbb{R}^{3}$, under ambient isotopy. To simplify the discussion, we will abuse notation and write $\gamma$ for the map as well as its image in $\mathbb{R}^{3}$. A link with $n$ components is a disjoint union of $n$ knots, $L=\gamma_{1},\cup\ldots\cup\gamma_{n}\subset\mathbb{R}^{3}$. We say two knots (or links) are equivalent if there is an orientation-preserving homeomorphism from $\mathbb{R}^{3}$ to $\mathbb{R}^{3}$ such that one knot (or link) is the image of the other.
For each $u\in\mathbb{S}^{2}$, the projection of a knot (or link), $\gamma$, in the direction $u$ provides a knot diagram (or link diagram) of $\gamma$ with crossings. In this paper we will assume diagrams with generic crossings.
A braid on $m$ strands is the disjoint union of $m$ intervals embedded in a solid cylinder, $B=B_{i}\colon I\to D^{2}\times I$, monotonically increasing with respect to $I$, such that $B_{i}(0)=(d_{i},0)$ and $B_{i}(1)=(d_{j},1)$ for each strand, so that the set of endpoints of strands is some permutation of the set of origins. Two braids are equivalent if there is an orientation-preserving homeomorphism from $D^{2}\times I$ to $D^{2}\times I$ such that one braid is the image of the other and at each time during the homeomorphism the image is also a braid. For each $u\in\mathbb{S}^{1}$, the projection of a braid in the direction $u$ provides a braid diagram of $B$ with crossings. In this paper we will assume braid diagrams with generic crossings, and we further use that each braid can be represented in a piecewise vertical diagram, that is the strands in the diagram are vertical expect in a (small) neighbourhood of a crossing. This result seems to be folklore; see for example “Reidemeister’s Theorem" and the discussion in [4].
Definition 2.1 (Closed braid).
A closed braid or braided link is the image of a braid under the map from the solid cylinder to the solid torus, $D^{2}\times I\to D^{2}\times\mathbb{S}^{1}{\subset\mathbb{R}^{2}\times\mathbb{C}\simeq\mathbb{R}^{4}}$, which sends $(x,y)\mapsto(x,e^{iy})$, where $e^{iy}$ gives the standard embedding from $\mathbb{R}/2\pi\mathbb{Z}$ into $\mathbb{C}\simeq\mathbb{R}^{2}$. Under the standard embedding $\mathcal{T}$ of the solid torus in $3$-dimensional Euclidean space, that is rotationally symmetric around the $z$-axis and orienting each strand positively the closed braid can be considered a link with $n\leq m$ components. Importantly, the orientation of each resulting component is aligned to a positive orientation on the core circle of the solid torus at all points. The braid index is the minimum number of strands required to form a closed braid equivalent to a given link.
We note that previous work has connected braids and braid groups with monodromy [8, 7, 24, 25], further motivating the connection which we explore in this paper.
The following result of Alexander will be essential to our result.
Theorem 2.2 (Alexander 1923 [2]).
Every knot or link is equivalent to the oriented image of a closed braid.
Of course this correspondence is not bijective as each link may be equivalent to many closed braids. An algorithmic alternative proof of this result was later given in [29]. They also showed that the number of elementary operations to obtain a braid diagram from a given link diagram with $n$ crossings and $p$ Seifert circles is at most $(p-1)(p-2)/2$ and the number of crossings in the resulting braid is at most $n+(p-1)(p-2)$. The braid index of a link is the smallest number of strands needed for a closed braid representation of the link. The braid index is equal to the minimal number of Seifert circles in any diagram of the braid [30]. We’ll now define the Seifert circles.
Definition 2.3 (Seifert Circle).
Given an oriented link diagram, by eliminating each crossing and connecting incoming strands with adjacent outgoing strands we obtain a diagram of oriented circles known as Seifert circles.
2.3 Vineyards
We make the following blanket assumption: We assume that all our persistence diagrams only contain a finite number of points. We note that this is always true for $C^{2}$ manifolds with a tame Morse functions, which is the setting we require. [^1] Note that we are not referred to tame knots but rather tame Morse functions in this work, as an embedded tame knot may have an infinite number of critical points in its distance or height filtration.
2.3.1 Persistence
We assume the reader is already familiar with homology, but we nevertheless give a crash course in persistence and refer the curious to [14, 21] for further details. Whereas classical homology is useful for studying topological features of data, it lacks the discernment to pick the correct scale to extract features from. Persistent homology remedies this shortcoming by studying nested sequences of sublevel sets called filtrations, thereby extracting features across many scales. Taking a manifold $\mathcal{M}$ and a nice function to filter it with, for example, the height function, yields the persistence module comprised of homology groups and linear maps induced by the inclusions between sublevel sets
$\ldots\to H(\mathcal{M}_{a_{i-1}})\to H(\mathcal{M}_{a_{i}})\to\ldots\to H(\mathcal{M}_{a_{j-1}})\to H(\mathcal{M}_{a_{j}})\to\ldots.$ (1)
Composing the maps between consecutive groups, we get a map between any two groups in the module. We say a homology class $\alpha\in H(\mathcal{M}_{a_{i}})$ is born at $\mathcal{M}_{a_{i}}$ if it is not in the image of the map from $H(\mathcal{M}_{a_{i-1}})$ to $H(\mathcal{M}_{a_{i}})$. If $\alpha$ is born at $\mathcal{M}_{a_{i}}$, it dies entering $\mathcal{M}_{a_{j}}$ if the image of the map from $H(\mathcal{M}_{a_{i-1}})$ to $H(\mathcal{M}_{a_{j-1}})$ does not contain the image of $\alpha$, but the image of the map from $H(\mathcal{M}_{a_{i-1}})$ to $H(M_{a_{j}})$ does. The persistence of $\alpha$ is the difference between the function values at its birth and its death. If the function is a Morse function on a manifold, then precisely one Betti number $\beta_{p}$ changes when the threshold passes a critical value. If the index of the corresponding critical point is $p$, then either a $p$-dimensional class is born, so $\beta_{p}$ increases by one, or a $(p-1)$-dimensional class dies, so $\beta_{p-1}$ decreases by one. We use a persistence diagram ${\rm Dgm}{({f})}$ to encode the birth-death information of all of the $p$-dimensional homology classes of $M$ arising from the sublevel set filtration induced by the filtering function $f$, where each birth/death pair becomes a point in $\mathbb{R}^{2}$. See Figure 4.
Figure 4: A torus ($\mathbb{S}^{l}\times\mathbb{S}^{1}$) embedded in $\mathbb{R}^{l+2}$. In the figure open black circles depict points of $H_{0}$- persistence, the red round dots depict points in ordinary $H_{l}$-persistence, the arrows indicate cycles that live forever (black for $H_{0}$, red for $H_{l}$, blue for $H_{1}$, and green for $H_{l+1}$).
2.3.2 Extended persistence
There are some drawbacks to the topological summary given by standard persistence, most notably points at infinity. For example, consider a sublevel set filtration of a shape embedded in $\mathbb{R}^{2}$ or $\mathbb{R}^{3}$ with nontrivial $H_{0}$ and $H_{1}$: at some point, connected components and loops are born, but never die, as they are present in all sublevel sets after their initial appearance. Even worse, if the input is non-generic and two $H_{1}$ cycles are born at the same height in the sublevel set filtration, then they give rise to identical persistence pairs.
To address this, Agarwal et. al. [1] established a pairing between all critical points of a height function on a 2-manifold, which Cohen-Steiner et. al. [10] extended to general manifolds with tame functions, leveraging Poincaré and Lefschetz duality to create a new sequence of homology groups where we begin and end with the trivial group. This guarantees that each homology class that is born will also die at a finite value, replacing all the problematic points paired with $\infty$, and guarantees a perfect matching on critical points. Let $H_{d}(\mathcal{M},\mathcal{M}^{a})$ denote the relative homology group of $\mathcal{M}$ with the superlevel set $\mathcal{M}^{a}=f^{-1}[a,\infty)$. Again assume we have a tame function and critical set $a_{1},\cdots,a_{k}$, and note that $H_{d}(\mathcal{M}_{a_{k}})=H_{d}(\mathcal{M})=H_{d}(\mathcal{M},\mathcal{M}^{a})$ for any $a>a_{k}$. From this, we can create a new sequence of homology groups
00 0 $\rightarrow H_{d}(\mathcal{M}_{a_{1}})\rightarrow H_{d}(\mathcal{M}_{a_{2}})\rightarrow\ldots\rightarrow H_{d}(\mathcal{M}_{a_{k-1}})$
$\rightarrow H_{d}(\mathcal{M}_{a_{k}})=H_{d}(\mathcal{M},\emptyset)\rightarrow\ H_{d}(\mathcal{M},\mathcal{M}^{a_{k}})$
$\rightarrow H_{d}(\mathcal{M},\mathcal{M}^{a_{k-1}})\rightarrow\ldots\rightarrow H_{d}(\mathcal{M},\mathcal{M}^{a_{1}})=H_{d}(\mathcal{M},\mathcal{M})=0,$ (2)
which we call the $d^{th}$ extended filtration sequence. We define the sequence $H_{d}(\mathcal{M}_{a_{1}})\rightarrow\ldots\rightarrow H_{d}(\mathcal{M}_{a_{k}})$ as the upwards sequence and the sequence $H_{d}(\mathcal{M},\mathcal{M}^{a_{k}})\rightarrow\ldots\rightarrow H_{d}(\mathcal{M},\mathcal{M}^{a_{1}})$ as the downwards sequence.
Note that this sequence fits the structure of a persistent module, so like standard persistence, it also has a unique interval decomposition. The only change is that we interpret the persistence points differently in this setting. Specifically, the points in the persistence diagram can be partitioned into three different groups: 1) the classes which are born and die in the upwards sequence, 2) the classes which are born and die in the downwards sequence, and 3) the classes which are born in the upwards sequence and die in the downwards sequence. Further, we associate the index of birth and death intervals with the value $a_{i}$ for both $H_{p}(\mathcal{M}_{a_{i}})$ in the the upwards sequence and $H_{p}(\mathcal{M},\mathcal{M}^{a_{i}})$ in the downwards sequence.
This creates three different classes of persistence pairs: those that correspond to a class that is born and dies in the upwards sweep (ordinary persistence points), those that correspond to a class that is born and then dies in the downwards sweep (relative persistence points), and those that correspond to a class that is born in the upwards sweep and dies in the downwards sweep (extended persistence points). We denote the class of ordinary, relative, and extended points as $\text{Ord}_{d}(f)$, $\text{Rel}_{d}(f)$, and $\text{Ext}_{d}(f)$, respectively. Note that the ordinary diagram is entirely above the diagonal; the relative diagram is entirely below the diagonal; while the extended persistence points can be on either side. See Figure 5.
Figure 5: A torus ($\mathbb{S}^{l}\times\mathbb{S}^{1}$) embedded in $\mathbb{R}^{l+2}$. In the figure open black circles depict points of ordinary or extended $H_{0}$- persistence, the red round dots depict points in ordinary $H_{l}$-persistence, the green squares (below the diagonal) indicate relative $H_{l+1}$-persistence and the blue crosses (below the diagonal) indicate relative $H_{1}$-persistence.
2.3.3 Vineyards
One can also study the persistence on a manifold $\mathcal{M}$ arising from multiple functions. As long as the functions are similar enough, the Stability Theorem [9] of persistent homology asserts that their associated persistence diagrams will also be similar; see also [28] for a discussion of the algebraic details of maps between “nearby" diagrams. More precisely, the bottleneck distance between two persistence diagrams of functions $g_{u},g_{v}\colon M\to{\mathbb{R}}$ is bounded from above by the infinity norm of their difference:
${W_{\infty}}(\mathscr{D}_{\cdot}(g_{u}),\mathscr{D}_{\cdot}(g_{v}))$ $\leq\|{g_{u}}-{g_{v}}\|_{\infty},$ (3)
in which $\mathscr{D}_{\cdot}(g)$ is the persistence diagram of $g$, a matching between the two diagrams is quantified by the supremum of the max-distances between matched points, and ${W_{\infty}}$ is the infimum over all possible matchings, in which we allow the introduction of points on the diagonal (where birth time is equal to death time), which we preferably match with points near to the diagnal to points on the diagonal. Intuitively, the bottleneck distance describes the worst disparity between the best matching of points in persistence diagrams: the worst disparity is the “bottleneck” preventing a smaller distance.
This gives rise to the concept of a vineyard [11, 28]. It formalizes the idea that a feature of $g_{u}$ is still recognizable in $g_{v}$, provided the two functions $u$ and $v$ are not too far apart. Features are points in the diagram, and the association is a matching between the points of $g_{u}$ and of $g_{v}$. Cohen-Steiner et al. [11] propose the so-called vineyard algorithm that traces the features (points) while continuously deforming $g_{u}$ into $g_{v}$, realized as an update to the reduced matrix arising from $g_{u}$. This update can be done in $O(n)$ time, as compared to the worst case $O(n^{3})$ time to compute and reduce the new matrix from scratch.
3 Contribution
In the context of persistence of the induced distance function, we need to introduce a number of conventions for monodromy to be well defined. In fact we will introduce monodromy both associated to an entire vineyard as well as to a single vine in the vineyard. We will then present our main results, demonstrating monodromy and braiding in detail through several examples.
3.1 Monodromy in vineyards: a geometric viewpoint
Let $\mathcal{M}$ be a manifold embedded in $\mathbb{R}^{d}$ and $\gamma:[0,2\pi]\to\mathbb{R}^{d}$ be a parametrization of a loop $\gamma$. Let $d(\cdot,\gamma(t)):\mathcal{M}\to{\mathbb{R}}$ be the function $x\mapsto d(x,\gamma(t))_{\mathcal{M}}$, where $d(x,\gamma(t))$ is the Euclidean distance from $x$ to $\gamma(t)$ and, as indicated, $d(x,\gamma(t))_{\mathcal{M}}$ is its restriction to the manifold $\mathcal{M}$. We will now identify the ends of the interval $[0,2\pi]$ – that is, we pass to $\mathbb{S}^{1}$. For each $t$ the function $d(\cdot,\gamma(t))_{\mathcal{M}}$ induces a filtration on $\mathcal{M}$, by the sub-level sets of the function. Therefore we can consider (for each $t$) the $l$th order persistence diagram of this filtration $\mathscr{D}_{l}(d(\cdot,\gamma(t))_{\mathcal{M}})$. The map
$\operatorname{CV}_{\mathcal{M}}:\mathbb{S}^{1}$ $\to\mathbb{S}^{1}\times\mathrm{Dgm}$
$t$ $\mapsto(t,\mathscr{D}_{l}(d(\cdot,\gamma(t))_{\mathcal{M}})),$
where $\mathrm{Dgm}$ is the space of persistence diagrams, is (trivially) a covering space of $\mathbb{S}^{1}$. We refer to $\operatorname{CV}_{\mathcal{M}}$ as the closed vineyard map. Thanks to [9] the points in the persistence diagram are (Lipschitz) continuous with respect to $t$. The map $\operatorname{CV}_{\mathcal{M}}$ is illustrated in Figure 6.
Figure 6: An illustration of the map $\operatorname{CV}_{\mathcal{M}}$. We indicate the persistence diagrams, which form the fibers, only explicitly in a number of places for visual clarity.
We will now define monodromy in two increasingly more complicated settings. Our definition is canonical only in the first setting, while the second definition involves a choice. We note that generically [28, page 3], there are no points of higher multiplicity in a persistence diagram and hence the connected components in the image of $\operatorname{CV}_{\mathcal{M}}$, that is the vines, are non-intersecting curves. We fix the assumption throughout this paper that our vineyards are generic in this sense.
Staying clear of the diagonal
This leads us to the definition of monodromy in the simplest setting (where we stay clear of the diagonal):
Definition 3.1.
Assume that there are no points of higher multiplicity in $\mathscr{D}_{l}(d(\cdot,\gamma(t))_{\mathcal{M}})$, for all $\gamma(t)$ and by extension the vines are non-intersecting curves. Assume moreover that all of the vines are disjoint from the diagonal.
Individual vine Given a point in the persistence diagram $\mathscr{D}_{l}(d(\cdot,\gamma(\tau))_{\mathcal{M}})$ for some $\tau\in\mathbb{S}^{1}$, write $\mathcal{V}(t)$ for the lift of $\gamma(t)$, which continuously assigns a point in $\mathscr{D}_{l}(d(\cdot,\gamma(t))_{\mathcal{M}})$ for every $t$ and yields the given point for $t=\tau$. We will now assume without loss of generality that $\tau=0$. We say that the vine $\mathcal{V}(t)$ exhibits monodromy if the start and end point of the lifted curve do not coincide. Similar to $\mathcal{V}(t)$, write $\tilde{\mathcal{V}}^{k}(t)$ for the analogous lift of $\gamma^{k}(t)$. Then $\mathcal{V}(t)$ exhibits monodromy of order $k$ if $k$ is the smallest integer strictly larger than 00, such that $\tilde{\mathcal{V}}^{k}(0)=\tilde{\mathcal{V}}^{k}(2\pi k)$. In other words the order of the monodromy is the number of points above any $t$ the connected component in the image of $\operatorname{CV}_{\mathcal{M}}$ that contains the given point in $\mathscr{D}_{l}(d(\cdot,\gamma(\tau))_{\mathcal{M}})$.
Vineyard Following an individual vine (with increasing time $t$) for a given point on $\mathscr{D}_{l}(d(\cdot,\gamma(0))_{\mathcal{M}})$ yields a point in $\mathscr{D}_{l}(d(\cdot,\gamma(2\pi))_{\mathcal{M}})=\mathscr{D}_{l}(d(\cdot,\gamma(0))_{\mathcal{M}})$. In other words the vines or vineyard induce a map $P_{\mathcal{V}}$ from $\mathscr{D}_{l}(d(\cdot,\gamma(0))_{\mathcal{M}})$ to itself, which permutes the points in the persistence diagram. We say that the vineyard has monodromy of order $k$ if $k$ is the order of the permutation, that is the smallest integer $k>0$ such that applying this permutation $k$ times yields the identity permutation.
Allowing to touch the diagonal
We next drop the assumption that the vines do not touch the diagonal, but still assume that none of the persistence diagrams contain points with higher multiplicity. To make sense of the other definition below and to provide an alternative formulation to the one above, we need the following observation: Let $\tilde{\mathcal{V}}_{\max}(t)$ be a vine defined on its maximal domain $(t_{\min},t_{\max})$, where $t_{\min},t_{\max}\in\mathbb{R}\cup\{\pm\infty\}$, that contains the point $\mathcal{V}(0)$. More formally $\tilde{\mathcal{V}}_{\max}(t)$ is the restriction to the maximal open interval such that the lift of the curve $\gamma:\mathbb{R}\to\mathbb{R}/2\pi\mathbb{Z}:t\mapsto t\;\textrm{mod}\;2\pi$, contains no limit points on the diagonal for this open interval.
If we are given a vine $\mathcal{V}(t)$ defined on $[0,2\pi]$ we call the vine $\tilde{\mathcal{V}}_{\max}(t)$ defined on its maximal domain that coincides with $\mathcal{V}(t)$ on the interval $[0,2\pi]$, the maximal extension of $\mathcal{V}(t)$. If the domain of a maximal extension is finite, then we say that the vine has a finite maximal extension.
If the number of points in each persistence diagram is finite (as we assume), then either both $t_{\min}$ and $t_{\max}$ are (plus/minus) infinite or neither of them are. This is clear because if one of them is infinite we must return to the same point in the persistence diagram after some time $2\pi k$ (since the number of points in the persistence diagram is finite), in which case $\tilde{\mathcal{V}}(t)$ is periodic and is defined on (,)($\infty$,$\infty$)( - ∞ , ∞ ).
Using this definition we have the following reformulation of the order of the monodromy in the setting of Definition 3.1:
Remark 3.2.
Assume that there are no points of higher multiplicity in $\mathscr{D}_{l}(d(\cdot,\gamma(t))_{\mathcal{M}})$, for all $\gamma(t)$ and by extension the vines are non-intersecting curves. Assume moreover that all of the vines are disjoint from the diagonal. The vine $\mathcal{V}(t)$ exhibits monodromy of order $k$ if and only if $k$ is the smallest positive integer such that $\tilde{\mathcal{V}}_{\max}(0)=\tilde{\mathcal{V}}_{\max}(2\pi k)$, where $\tilde{\mathcal{V}}_{\max}(t)$ is the maximal extension of $\mathcal{V}(t)$.
Figure 7: A schematic overview of how we find monodromy in the vineyard. The gray part of the knot does not contribute to monodromy, and is not shown in the vineyard.
In this setting, that is where we allow vines to contain points in the diagonal (as limit points), we make the following definition:
Definition 3.3.
Assume that there are no points of higher multiplicity in $\mathscr{D}_{l}(d(\cdot,\gamma(t))_{\mathcal{M}})$, for all $\gamma(t)$ and by extension the vines are non-intersecting curves. We’ll follow the same notation as above.
Individual vine Let $\mathcal{V}(t)$ be the vine and $\tilde{\mathcal{V}}_{\max}(t)$ is its maximal extension. If its maximal domain is (,)($\infty$,$\infty$)( - ∞ , ∞ ), then the vine does not touch the diagonal and Definition 3.1 and the reformulation in Remark 3.2 both apply. Let us now assume that $t_{\min},t_{\max}$ are finite. By reparametrizing we can assume (without loss of generality) that $t_{\min}=0$. The order of monodromy is now defined as the smallest positive integer $k$, such that $t_{\max}<2\pi k$. As before, we’ll say that the monodromy is trivial if $k=1$.
We’ll also define the completion along the diagonal of a vine $\mathcal{V}$. Let $\tilde{\mathcal{V}}_{\max}:(t_{\min},t_{\max})\to\mathscr{D}_{l}(d(\cdot,\gamma(t))_{\mathcal{M}})$ be the maximal extension of $\mathcal{V}$, and write
$v_{\min}=\lim_{t\searrow t_{\min}}\tilde{\mathcal{V}}_{\max}(t)$
$v_{\max}=\lim_{t\nearrow t_{\max}}\tilde{\mathcal{V}}_{\max}(t)$
for the two limit points of the vine on the diagonal. Let $k$ be the order of monodromy of $\mathcal{V}$ we define the completing diagonal vine $\mathcal{V}_{D}$ as
$\mathcal{V}_{D}:$ $(t_{\max},t_{\min}+2\pi k)\to\mathscr{D}_{l}(d(\cdot,\gamma(t))_{\mathcal{M}})$
$t\mapsto\left(t,v_{\max}\left(1-\frac{t-t_{\max}}{t_{\min}-t_{\max}+2\pi k}\right)+v_{\min}\frac{t-t_{\max}}{t_{\min}-t_{\max}+2\pi k}\right).$
The start and end points of the concatenation $\mathcal{V}_{D}\circ\tilde{\mathcal{V}}_{\max}$ coincide and therefore by identifying $t_{\min}$ and $t_{\min}+2\pi k$ we can consider this to be a map on $\mathbb{R}/2\pi k\mathbb{Z}$. The set $\mathbb{R}/2\pi k\mathbb{Z}$ can be viewed as $k$ cover of the circle $\mathbb{R}/2\pi\mathbb{Z}$, where we identify $\mathbb{R}/2\pi\mathbb{Z}$ with the loop $\gamma$. Composing with this cover map gives a map $\mathcal{V}_{C}:\gamma(t)\mapsto\mathscr{D}_{l}(d(\cdot,\gamma(t))_{\mathcal{M}})$, which exhibits monodromy in the way we defined above. We call $\mathcal{V}_{C}$ the completion of the vine $\mathcal{V}$.
Vineyard We now call the vines whose maximal extension do not intersect the diagonal (or equivalently those whose maximal domain is (,)($\infty$,$\infty$)( - ∞ , ∞ )) non-rooted vines, while we call those that intersect the diagonal (or equivalently those whose maximal domain has finite length) rooted vines. We write $k_{1},\dots,k_{n}$ for the orders of monodromy of non-rooted vines and write $l_{1},\dots,l_{m}$ for the orders of monodromy of the rooted vines. Let $l_{\max}=\max\{l_{1},\dots,l_{m}\}$, then we define the order of monodromy of the vineyard as the smallest $k$ that is a common multiple of $k_{1},\dots,k_{n}$ and is larger than $l_{\max}$. We also call $k$ the common order of monodromy of all the vines.
We can complete the vineyard in the same way as before, but with $k$ the common order of monodromy of all the vines, instead of an individual vine. We call the result the completed vineyard.
We stress that there may be self-intersections on the diagonal of the completed vineyard. We note again here that if the vineyards are non-generic, the situation is significantly more complex as vines can intersect [28]; we will not consider the non-generic case any further in this work.
3.2 Braids and Morse theory
We say that a closed braid $B$ is $(\epsilon,R)$-embedded in $\mathbb{R}^{3}$ if the closed braid is contained in the $\epsilon$-thickening of the circle $C_{h}(0,R)$ of radius $R$ contained in the horizontal plane. We will tacitly assume that $\epsilon\leq R$.
We say that the maximal angle that a strand of a braid makes with the horizontal direction is the maximal braid angle $\theta_{B}$. For a closed braid we say that the maximal braid angle $\theta_{B}$ is the angle that the tangent line to a strand at a point $p$ in the braid makes with the normal of the plain $P$ defined by $p$ and the $z$-axis.
We can consider a small [^2] The offset should be small enough such that no self-intersection, nor intersections with the $z$-axis occur. $(l+1)$-dimensional $\alpha$-offset $\mathcal{M}$ of a closed braid $B$. This offset is defined as follows: For $l=0$, we do not add an offset. For $l\geq 1$, we consider the embedding $\mathcal{T}\times 0$ in $\mathbb{R}^{3}\times\mathbb{R}^{l-1}$, where $\mathbb{R}^{3}$ is the space that contains the standard embedding of the torus, in the sense of Definition 2.1. We take $\mathcal{M}$ to be the offset of the braid $B\times 0$ in $\mathbb{R}^{3}\times\mathbb{R}^{l-1}$, that is the boundary of $(\mathcal{T}\times 0)\oplus B(0,\alpha)$, where $\oplus$ denotes the Minkowski sum. The resulting manifold $\mathcal{M}$ can then be embedded in $\mathbb{R}^{d}$ where the $\mathbb{R}^{3}$ that contained $\mathcal{T}$ corresponds to the first three coordinates. The maximal braid angle in this context is defined as follows: Let $p\in\mathcal{M}$ and $T_{p}\mathcal{M}$ be its tangent space, then the maximal braid angle is the angle between the normal $n$ of the hyperplane $P\times\mathbb{R}^{d-3}$, where $P$ is the plane defined by the $z$-direction and the point $p$.
Lemma 3.4.
Let $B$ be $(\epsilon,R)$-embedded closed braid and suppose that $\mathcal{M}$ is its $(l+1)$-dimensional $\alpha$-offset, with braid angle $\theta_{B}$. If $p\in\mathbb{R}^{d}$ satisfies $d(p,C_{h}(0,R))\leq\eta$, $C_{h}(0,R)$ is parametrized by $s(\theta)$, and the closest point $p^{\prime}$ of $p$ on $C_{h}(0,R)$ satisfies $p^{\prime}=s(0)$. Then the function $x\mapsto d(x,p)_{\mathcal{M}}$ has no critical points at the closest point $\beta(\theta)$ on $B$ to $x$ to as long as
$\frac{\theta}{2}+\theta_{B}+\arcsin\frac{\epsilon}{2R\sin(\frac{\theta}{2})-\eta}+\arcsin\frac{\eta}{2R\sin(\frac{\theta}{2})}<\frac{\pi}{2}.$ (4)
This implies in particular that there is no topological change as long as (4) is satisfied.
We refer to Figure 10 for a situation sketch.
The proof of this lemma depends on one of the Morse theorems, see e.g. [20, Theorem 3.1]:
Theorem 3.5.
Suppose $f$ is a smooth real-valued function on $\mathcal{M}$, $a<b$, $f^{-1}[a,b]$ is compact, and there are no critical values between $a$ and $b$. Then $M^{a}$ a is diffeomorphic to $M^{b}$, and $M^{b}$ deformation retracts onto $M^{a}$.
Proof 3.6 (of Lemma 3.4).
We write $\beta(\theta)$ for a parametrization of $B$ according to the angle of the circle $C_{h}(0,R)$. We parametrize the circle as $s(\theta)=(R\sin\theta,R\cos\theta,0,\dots,0)$. Using this notation we note that the gradient of $d(y,p)_{\mathcal{M}}$ is zero at $y=x$ if for its closest point on $B$, that is $\beta(\theta)$, we have that $\langle{\beta}^{\prime},\beta-p\rangle=0$, where $\beta^{\prime}$ denotes the derivative of $\beta$ with respect to $\theta$. This is equivalent to
$\angle{\beta}^{\prime},\beta-p=\pi/2.$
By assumption we have that
$\angle{\beta}^{\prime},s^{\prime}\leq\theta_{B}.$
Moreover, because $|\beta(\theta)-s(\theta)|\leq\epsilon$, we also have
$\sin(\angle\beta-p,s-p)\leq\frac{\epsilon}{|s-p|},$
as can be seen from Figure 8.
Figure 8: The angle estimate for $\angle\beta-p,s-p$.
Using the triangle inequality of angles (or points on the sphere) we find that
$|\angle({\beta}^{\prime},\beta-p)-\angle({s}^{\prime},s-p)|\leq\theta_{B}+\arcsin\frac{\epsilon}{|s-p|}.$ (5)
Let us now write $d=|s-p|$, $p^{\prime}$ for the closest point projection of $p$ on $C_{h}(0,R)$ and $\omega=\angle({s}^{\prime},s-p^{\prime})$. By reparameterization we can assume that $p^{\prime}=s(0)$. With this assumption we see by the construction in Figure 9 that $\omega=\frac{\theta}{2}$. Moreover we have that $|p^{\prime}-s|=2R\sin(\frac{\theta}{2})$
Figure 9: The construction for the angle $\omega=\angle({s}^{\prime},s-p^{\prime})$.
By the same argument as given in Figure 8 we have that
$\sin\angle(s-p^{\prime},s-p)\leq\frac{\eta}{|s-p^{\prime}|},$
so that together with (5) we find that
$\left|\angle({\beta}^{\prime},\beta-p)-\frac{\theta}{2}\right|$ $\leq\theta_{B}+\arcsin\frac{\epsilon}{|s-p|}+\arcsin\frac{\eta}{|s-p^{\prime}|}$
$\leq\theta_{B}+\arcsin\frac{\epsilon}{|s-p^{\prime}|-\eta}+\arcsin\frac{\eta}{|s-p^{\prime}|}$ (by the triangle inequality)
$\leq\theta_{B}+\arcsin\frac{\epsilon}{2R\sin(\frac{\theta}{2})-\eta}+\arcsin\frac{\eta}{2R\sin(\frac{\theta}{2})}$
This means that if
$\frac{\theta}{2}+\theta_{B}+\arcsin\frac{\epsilon}{2R\sin(\frac{\theta}{2})-\eta}+\arcsin\frac{\eta}{2R\sin(\frac{\theta}{2})}<\frac{\pi}{2}$
then $\angle({\beta}^{\prime},\beta-p)\neq\pi/2$ and hence there are no critical points. The fact that there is no topological change follows from Theorem 3.5.
Remark 3.7.
The important conclusion from the bound of Lemma 3.4 is that by choosing $\epsilon<\eta\simeq\theta_{B}$ small compared to $\min\{R,\pi\}$, the bound (4) is satisfied as long as $\theta_{B}\ll\theta<\pi-4\theta_{B}$.
We say that an oriented closed braid $B$ that is $(\epsilon,R)$-embedded is $\delta$-circular if its parametrization according to arc length $\beta$ satisfies $\left|\frac{\beta(t)}{R^{2}}+\ddot{\beta}(t)\right|<\delta$, where we use Newton’s notation for the derivative.
Lemma 3.8.
Let $B$ be an oriented $(\epsilon,R)$-embedded, $\delta$-circular closed braid, with braid angle $\theta_{B}$. If $p\in\mathbb{R}^{d}$ satisfies $d(p,C_{h}(0,R))\leq\eta$, $C_{h}(0,R)$ is parametrized by $s(\theta)$, and the closest point $p^{\prime}$ of $p$ on $C_{h}(0,R)$ satisfies $p^{\prime}=s(0)$.
If an oriented $(\epsilon,R)$-embedded, $\delta$-circular closed braid with $n$ strands, with braid angle $\theta_{B}$, $\epsilon\ll R$, and
$\frac{6\epsilon}{R}+\delta(R+\eta)\leq\frac{1}{R^{2}}(R-\epsilon)(R-\eta),$
and $\epsilon<\eta\simeq\theta_{B}$ are small compared to $\min\{R,\pi\}$, then $d(\cdot,p)|_{B}$ has $2n$ critical points, $n$ maxima and $n$ minima. Let $l\geq 1$. If $\mathcal{M}$ is a $(l+1)$-dimensional $\alpha$ offset $\mathcal{M}$ of the same type of braid satisfying the same conditions then $d(\cdot,p)|_{\mathcal{M}}$ has $4n$ critical points, $n$ maxima and $n$ minima and $2n$ saddle points of index $l$.
Proof 3.9.
Because the conditions of Lemma 3.4 are satisfied we know that there are no Morse critical points unless $\theta\simeq 0$ or $\theta\simeq\pi$ so we focus on establishing that there is only one critical point per strand at $\theta\simeq 0$ or $\theta\simeq\pi$ for $B$ and or two per strand at $\theta\simeq 0$ or $\theta\simeq\pi$ in the case of $\mathcal{M}$.
The proof for the braid case is the difficult step, and we will see below that the statement for $\mathcal{M}$ follows immediately. The idea of the proof is the following: Because the closed braid is $\delta$-circular its parametrization $\beta$ satisfies $\left|\frac{\beta(t)}{R^{2}}+\ddot{\beta}(t)\right|<\delta$. This means that $\beta$ is forced to turn inward towards the centre of $C_{h}(0,R)$, following the circle $C_{h}(0,R)$, where 00 is the origin of Euclidean space. This in turn implies that $\beta$ cannot satisfy $\frac{d}{dt}|\beta-p|^{2}(t)=0$ for two times $t$ that are relatively close. In other words the curve cannot be tangent to some sphere (not necessarily of the same radius) centred at $p$.
The way that we establish that $\frac{d}{dt}|\beta-p|^{2}(t)$ cannot be zero for two nearby values is by establishing that if $\frac{d}{dt}|\beta-p|^{2}(t)=0$ then
$\frac{d}{dt}\left(\frac{d}{dt}|\beta-p|^{2}\right)(t)=\frac{d^{2}}{dt^{2}}|\beta-p|^{2}(t)$
is large. If we write $\frac{\beta(t)}{R^{2}}+\ddot{\beta}(t)=\Delta(t)$ and suppress $t$ from the notation, we see that
$\frac{d^{2}}{dt^{2}}|\beta-p|^{2}(t)$ $=2\langle\ddot{\beta},\beta-p\rangle+2\langle\dot{\beta},\dot{\beta}\rangle$
$=2\langle\ddot{\beta},\beta-p\rangle+2$ (because $|\dot{\beta}|=1$)
$=2\left\langle-\frac{\beta}{R^{2}}+\Delta,\beta-p\right\rangle+2$
$=-2\left\langle\frac{\beta}{R^{2}},\beta\right\rangle+2\left\langle\Delta,\beta-p\right\rangle+2+2\left\langle\frac{\beta}{R^{2}},p\right\rangle$ (6)
We can now examine the first three terms in (6):
  • Because $B$ is $(\epsilon,R)$-embedded $R-\epsilon\leq|\beta|\leq R+\epsilon$, so that
    $-2\left(1+\frac{\epsilon}{R}\right)^{2}\leq-2\left\langle\frac{\beta}{R^{2}},\beta\right\rangle\leq-2\left(1-\frac{\epsilon}{R}\right)^{2},$
    which, if $\epsilon\ll R$, simplifies to
    $-2-\frac{6\epsilon}{R}\leq-2\left\langle\frac{\beta}{R^{2}},\beta\right\rangle\leq-2+\frac{6\epsilon}{R}.$
  • Because $|\Delta(t)|=|\frac{\beta(t)}{R^{2}}+\ddot{\beta}(t)|\leq\delta$, Cauchy-Schwarz yields that $2|\left\langle\Delta,\beta-p\right\rangle|\leq\delta|\beta-p|\leq\delta(R+\eta)$.
This implies that these three terms are close to zero, i.e.
$\frac{d^{2}}{dt^{2}}|\beta-p|^{2}(t)\simeq 2\left\langle\frac{\beta}{R^{2}},p\right\rangle.$
Because $\left|2\left\langle\frac{\beta}{R^{2}},p\right\rangle\right|$ is lower bounded by $\frac{1}{R^{2}}(R-\epsilon)(R-\eta)$ assuming that the angle between $\beta$ and $p$ is no more than $45$ degrees (or more than $135$), the first part of the result now follows if, the angle between $\beta$ and $p$ is no more than $45$ degrees (or more than $135$), $\epsilon\ll R$,
$\frac{6\epsilon}{R}+\delta(R+\eta)\leq\frac{1}{R^{2}}(R-\epsilon)(R-\rho_{\max}).$
For the second part, note that there is a one-to-one correspondence between pairs of critical points of the distance function to a fixed point on an offset and the critical points of the distance function to the same fixed point on the curve $\beta$ itself.
Given these $4n$ critical points, we can now consider the persistence diagram that results. Although we work with extended persistence throughout this paper in order to avoid points at infinity and establish a perfect pairing of critical points, we note that in fact for the purposes of establishing monodromy in our knot offset, it suffices to restrict our attention to the behavior of the ordinary points as well as one single extended point, all above the diagonal, as those that contribute to the construction of knot or link as mentioned in Theorem 1.2 are born and die ‘on the way up’, that is in the first line of the sequence in (2).
Corollary 3.10.
Under the same assumptions as in Lemma 3.8, we have that the maxima and minima of $d(\cdot,p)|_{B}$ and $d(\cdot,p)|_{\mathcal{M}}$ as well as the saddle points of $d(\cdot,p)|_{\mathcal{M}}$ can be divided into two groups, one group occurring at low values and corresponding to $H_{0}$ births of cycles in the persistence diagram and one group occurring at high values and corresponding to $H_{0}$ deaths in the persistence diagram. For $H_{l}$ the situation is identical except for one change, namely that a single birth occurs at a high value. In ‘ordinary’ persistence theory this cycle lives forever, while in extended persistence it dies at the global minimum, and in fact lies below the diagonal.
Proof 3.11.
The only thing in this corollary that requires an extra argument on top of Lemma 3.8, is the correspondence of the critical points with the births and deaths respectively: Because, by Remark 3.7 there are no Morse critical points unless $\theta\simeq 0$ or $\theta\simeq\pi$, we know that $B(p,r)\cap B$ (respectively $B(p,r)\cap\mathcal{M}$) with $r\simeq R$ consists of $n$ topological line segments (topological cylinder segments $\mathbb{S}^{l+1}\times[0,1]$). While at for $r>2R+\eta+\epsilon$ $B(p,r)\cap B$ (respectively $B(p,r)\cap\mathcal{M}$) consists of a topological circle or knot (its offset respectively). See Figure 10. The only way that we can achieve this with the number of critical points we found in Lemma 3.8 is if the births and deaths occur as described in the statement of the corollary.
Figure 10: A figure illustrating the statement of Corollary 3.10, highlighting the intersection of ouroboros (in yellow) and family of 3d growing spheres that highlight the $n$ births and $n$ deaths in $H_{0}$. See also Figure 11 for a different view of the level sets.
Let $f_{1}$ and $f_{2}$ be two functions on the same manifold $\mathcal{M}$. We say that these two functions are handle-equivalent if the handle decomposition is the same and the times of insertion of these handles are also identical. We have the following observation.
Corollary 3.12.
Let $B$ be $(\epsilon,R)$-embedded closed braid and suppose that $\mathcal{M}$ is its $(l+1)$-dimensional $\alpha$-offset, with braid angle $\theta_{B}$. We have that the $B$ is a circle and the manifold $\mathcal{M}$ is diffeomorphic to the torus $\mathbb{S}^{l+1}\times\mathbb{S}^{1}$. If $B$ is a braid with $n$ strands and let $p\in O_{\rho_{\min},\eta,\Psi}$, then $d(\cdot,p)|_{B}$ and $d(\cdot,p)|_{\mathcal{M}}$ are handle equivalent to the height function of the embedding depicted in Figure 11.
Figure 11: Top: The closed braid. Bottom: The embedding with equivalent height function. The arrow indicates the direction of the height function. The $b^{i}$s indicate the birth times and $D^{J}$, where $J$ is a roman numeral, the death times. We stress that the picture should be interpreted in a 3D way, and in particular $D^{II}$ does not have to be larger than $D^{J}$, with $J\neq II$. We stress that critical point with the highest value of the Morse function ($D^{II}$ in the figure) corresponds to a death only in extended persistence, in non-extended persistence, only a $1$-cycle is born there.
Remark 3.13.
We stress that given a closed braid $B$ it is not difficult to adjust the embedding such that it is $(\epsilon,R)$-embedded and $\delta$-circular with $\epsilon$ and $\delta$ as small and $R$ as large as you like. Because if $B$ is a closed braid, its parametrization $\beta(t)$ can be written as $\beta(\theta)=n(\theta)+\rho(\sin(\theta),\cos(\theta),0,\dots,0)$ for some $\rho>0$, with $n$ normal to $(\cos(\theta),-\sin(\theta),0,\dots,0)$. By redefining $\beta(\theta)=\tilde{\epsilon}n(\theta)+R(\sin(\theta),\cos(\theta),0,\dots,0)$ for sufficiently small $\tilde{\epsilon}$ the $\delta$-circular $(\epsilon,R)$-embedding can be achieved. By the same argument the observation loop (to be defined in the proof of Theorem 1.2) can be made arbitrarily close in terms of $\eta$ to a circle with radius $R$, so that it has reach as close to $R$ as one likes, by [17, Theorem 4.19], and hence has a nice tubular neighbourhood of that size.
3.3 Vineyard braiding
We will now prove (one of) the main statement(s) of the paper:
See 1.2
Proof 3.14.
The proof of this theorem is constructive. Thanks to Alexander’s theorem, Theorem 2.2, a given link can be represented as a closed braid. To every connected component of the closed braid we add an extra loop (using a Reidemeister move of type I) as in Figure 1. This Reidemeister move may introduce extra crossings, but at most $\mathcal{O}(s\cdot K)$, where $s$ denotes the number of strands in the original closed braid and $K$ the number of connected components of the loop. We now write $C$ for the number of crossings of the resulting closed braid. We write $n$ for the resulting number of strands, for which we see that $n=s+K$. We define/construct a particular embedding of the closed braid such that the original braid appears in the vineyard. Essentially, the closed braid diagram is parametrized by $\rho$ and $\theta$, while the embedding of the closed braid in $\mathbb{R}^{3}$ is parametrized by $\rho$, $\theta$, and $h$, as in Figure 13. Roughly speaking the parameter $\rho$ corresponds to the birth time, $\theta$ is the parameter of the observation loop $\gamma$ (to be defined in Step 3) and $h$ corresponds to the death time, again see Figure 13.
Our construction proceeds stepwise:
  • STEP 1 We start with an embedding which is close to its braid annular diagram, by which we mean that the embedding of the closed braid lies in a neighbourhood of an annulus in the plane and the braid is planar with the exception of small neighbourhoods of the points of crossing. We write $\theta,r$ for the coordinates of the annulus, which are polar coordinates in the plane restricted to the annulus.
  • STEP 2 We then modify (if necessary) the braid such that the crossings are equally parsed on one side of the annulus, that is, if $\theta$ is the angle that parametrizes the annulus, see Figure 12, then the crossings are contained in the interval $[0+\frac{\pi}{8(C+1)},\pi-\frac{\pi}{8(C+1)}]$, where $C$ is the number of crossings. By equally parsed we mean that there is only one crossing in each of the intervals $[\pi\frac{\tilde{j}}{8(C+1)},\pi\frac{\tilde{j}+1}{8(C+1)}]$ where $\tilde{j}\in\{8,16,\dots,8C\}$.
  • STEP 3a We now modify the embedding of the braid in an angular interval $[\pi-\frac{\pi}{4(C+1)},2\pi]\cup[0,\frac{\pi}{2(C+1)}]$. This interval should be interpreted in a periodic manner. We do so by twisting the annulus (and by extension the almost annular braid) $90$ degrees in the direction orthogonal to the plane into which the annulus was originally embedded, see Figure 1. We do so in such a way that the twisted annulus and by extension part of the braid in the angular interval $[\pi+\frac{\pi}{8(C+1)},2\pi-\frac{\pi}{8(C+1)}]$ is now close to a cylinder. We do this is a way that preserves cylindrical coordinates, that is, if $\theta$ was the planar angular coordinate of a point on the annulus, after twisting $\theta$ is the cylindrical coordinate of the corresponding point. We denote the resulting twisted annulus by $\mathcal{A}_{T}$.
  • STEP 3b We define an observation loop $\gamma$ to be the curve that follows that twisted annulus on the outside at a constant distance (less than $\eta$, with $\eta$ as in Lemmas 3.4 and 3.8), see e.g. Figure 2. We also define coordinates $\tilde{\theta},\rho,\tilde{\psi}$ with respect to this observation loop in a tubular neighbourhood of size $\mathcal{O}(\eta)$ of the observation loop. This is possible thanks to Remark 3.13. The coordinate $\rho$ of a point $y$ in the tubular neighbourhood is the distance to $\gamma$. The coordinate $\tilde{\psi}$ is the angle between a point $y$ in the tubular neighbourhood of the curve, its closest point on curve $\pi_{\gamma}(y)$, and the closest point of this point on the annulus $\pi_{\mathcal{A}_{T}}(\pi_{\gamma}(y))$. See Figure 12.
    Figure 12: Figure illustrating the notation (the coordinates $(\theta,\rho,\tilde{\psi})$) in step 3b of Theorem 1.2. The observation loop is the outer loop shown in dark blue, and (the flat part of) the twisted annulus is shown in grey.
Intermezzo: The persistence diagram for a point on the observation curve
Before we continue with the construction we now discuss the persistence diagram for a given point. In the following step we’ll further modify the closed braid in the angular interval $[\pi-\frac{\pi}{4(C+1)},2\pi]\cup[0,\frac{\pi}{2(C+1)}]$, but for now we consider the braid fixed. We also assume for now that the link has only one connected component, that is it is a knot. Thanks to Lemmas 3.4 and Corollaries 3.10-3.12 we know that the braid and observation curve are chosen in such a way that all the births occur first and then all the deaths occur, in both distance order as well as in order along the braid. Let us in particular consider the equivalent embedding of Corollary 3.12, where we let $b^{1},b^{2},\dots b^{n}$ be the births as they occur in order following along the braid $B$, as in Figure 11; note that we are slightly abusing notation here, as we are identifying the births with the Morse critical points. Next, let $b^{1}_{0},b^{2}_{0},\dots,b^{n}_{0}$ respectively $b^{1}_{l},b^{2}_{l},\dots,b^{n}_{l}$, where the subscript indicates if the birth occurs in 00 or $l$-homology, be the corresponding ordered births of $\mathcal{M}$. Note that we always have $n$ births unless $l=1$, when instead there are $n+1$ births. Again we emphasize that the births are ordered they occur following $B$ starting with the first birth, not in order of birth time according to the distance filtration; See Figure 11. Similarly, let $D^{I},D^{II},\dots,D^{N}$ be the deaths as they occur in order if we consider $B$, and $D^{I}_{0},D^{II}_{0},\dots,D^{N}_{0}$ respectively $D^{I}_{l},D^{II}_{l},\dots,D^{N}_{l}$ be the deaths if we consider them on the offset of $B$, $\mathcal{M}$. Again we assume that the deaths are ordered as they occur following $B$, not in order of deaths time.
We will assume without loss of generality that $b^{1}$ is the lowest birth value. Because of the elder rule the death of the cycle created at $b^{1}$ dies (in extended persistence) at the maximal death value, that is $\max_{J}D^{J}$, for $B$. Similarly $b^{1}_{0},b^{1}_{l}$ respectively die at $\max_{J}D^{J}_{0}$, $\max_{J}D^{J}_{l}$ respectively for $\mathcal{M}$.
The other persistence points are less straightforward, as they follow the mergers of the sublevel sets of the equivalent embedding of Corollary 3.12; see Figure 11. However, the most important case for us will be the following: Consider first any ordinary $H_{0}$ points, so we have $b^{1},b^{j},b^{k}$ with $1\neq j<k\neq 1$. Assume that $b^{1}$ is the earliest birth. Suppose that the death times are all larger than all birth times and are ordered as follows
$\max\{D^{I},D^{II},\dots,D^{J-2},D^{K+1},D^{K+2},\dots D^{N}\}$ $<\min\{D^{J-1},D^{K}\}$
$\max\{D^{J-1},D^{K}\}$ $<\min\{D^{J},\dots,D^{K-1}\},$ (7)
then the connected component born at time $b^{j}$ merges with the connected component born at time $b^{1}$ at time $D^{J-1}$ and the connected component born at time $b^{k}$ merges with the connected component born at time $b^{1}$ at time $D^{K}$ for $B$. In other words under these conditions there are points $(b^{j},D^{J-1})$ and $(b^{k},D^{K})$ in the persistence diagram.
We now consider $l$-homology persistent homology of the manifold $\mathcal{M}$. In general, as in Figure 5, the births in $l$-homology follow the births in 00-homology closely. By this we mean the following: As before we write $B$ for the braid and $\mathcal{M}$ for its $(l+1)$-offset, and we compare the same Morse function (distance to a point, or height if we consider the equivalent embedding) on the these two spaces ($B$ and $\mathcal{M}$, respectively). The birth of 00-cycles ($b^{1},b^{2},b^{3}$ in Figure 11) on the braid $B$ are close both geometrically and with regard to the value of the Morse function (the distance to a point or the height function, where in the latter case we consider the equivalent embedding) to the critical points that give rise to the birth of 00-cycles and $l$-cycles on $\mathcal{M}$. There is also a simple correspondence with one exception between deaths of 00-cycles for the braid and deaths of 00- and $l$-cycles on $\mathcal{M}$, meaning that the critical points that correspond to deaths for $B$ are close to a pair of critical points (one maximum and one saddle) that correspond to deaths in 00- and $l$-homology. The exception is the the last death of a 00-cycle in extended persistence on $B$; here we instead have a Morse critical point which corresponds to the birth of a $1$-cycle for the braid and which lives forever in the non-extended persistence, but corresponds to a death of a 00-cycle in extended persistence. There are again two corresponding nearby critical points on $\mathcal{M}$ for this final critical point on $B$, however in this case the saddle corresponds to the birth of a $1$-cycle (which corresponds to $\mathbb{S}^{1}$ in $\mathbb{S}^{1}\times\mathbb{S}^{l}$). This means that there is an extra point in the $l$-persistence diagram if $l=1$. Most importantly this point can be distinguished by the fact that its birth time is much higher than all other points in the persistence diagram. However, because the births and deaths of $B$ and $\mathcal{M}$ are so intimately linked except for the final death, we have the following: Consider $b^{1}_{l},b^{j}_{l},b^{k}_{l}$ and $1\neq j<k\neq 1$. Assume that $b^{1}$ is the earliest birth. If $l=1$ further assume that $b^{m}_{l=1}$ is the final birth and $j\leq m\leq k-1$. If now moreover the death times are ordered as follows
$\max\{D^{I}_{l},D^{II}_{l},\dots,D^{J-2}_{l},D^{K+1}_{l},D^{K+2}_{l},\dots D^{N}_{l}\}$ $<\min\{D^{J-1}_{l},D^{K}_{l}\}$
$\max\{D^{J-1}_{l},D^{K}_{l}\}$ $<\min\{D^{J}_{l},\dots,D^{K-1}_{l}\},$ (8)
then the $l$-cycle born at time $b^{j}_{l}$ merges with the $l$-cycle born at time $b^{1}_{l}$ at time $D^{J-1}_{l}$ and the $l$-cycle born at time $b^{k}_{l}$ merges with the $l$-cycle born at time $b^{1}_{l}$ at time $D^{K}_{l}$ for $\mathcal{M}$. In other words under these conditions there are points $(b^{j}_{l},D^{J-1}_{l})$ and $(b^{k}_{l},D^{K}_{l})$ in the persistence diagram.
  • STEP 4 As discussed in the intermezzo the first birth is always coupled to the last death in the persistence diagram. We write $b^{j,c}_{k}(t)$ and $D^{J,c}_{k}(t)$ for the births, deaths respectively of $d(\cdot,\gamma(t))|_{\mathcal{M}}$ per connected component $c$, where we stick to the convention that $b^{1,c}_{k}(t)$ is the first birth (in $k$-homology) for each connected component $c$. We use similar notation for $B$. Here, we note that although this distance value is continuous, the Morse critical point where this minimum is attained at is not continuous. A similar effect was called a Faustian interchange in [26]. With this notation we can conclude that our first observation implies that for each $c$ there is a point $(b^{1,c}_{k}(t),\max_{J}D^{J,c}_{k}(t))$ in the vineyard at level $t$ and the vine consisting of these points if closed (by identifying the vineyard at times 00 and $2\pi$) will yield a circle for each $c$. To put it differently this will lead to a surgery as depicted in Figure 13. Finally, we also note that if $l=1$ there is an additional $1$-cycle, that is born much later than all the other $1$-cycles.
    Figure 13: On the left we see a particular closed braid that we call the Ouroboros. On the right we see the $\theta$ and $\rho$ coordinates in the vineyard, where the $h$ coordinate is in the direction orthogonal to the plane. Here we identify $\rho$ with $R-b^{j}$ and $h$ with $D^{J^{\prime}}-2R$, where $D^{J^{\prime}}$ is the death time of the cycle born at $b^{j}$, corresponding to a strand as indicated in the figure. The elder rule induces surgery indicated with a dashed line (as depicted on the right).
    We further note that because vines are continuous (and even Lipschitz) thanks to [9], the only thing which we need to worry about is the crossings, because if all the birth values are distinct the birth values give precisely the $\rho$ coordinates in Figure 13, death times corresponding to the coordinate $h$ do not matter.
    We recall that there is only one crossing per (angular) interval $[\pi\frac{\tilde{j}}{8(C+1)},\pi\frac{\tilde{j}+1}{8(C+1)}]$ where $\tilde{j}\in\{8,16,\dots,8C\}$ and no crossings in $[\pi+\frac{\pi}{8(C+1)},2\pi-\frac{\pi}{8(C+1)}]$. Because the death times only matter for the crossing we can change the death times between crossings without creating topological problems. For a crossing in the interval $[\pi\frac{\tilde{j}}{8(C+1)},\pi\frac{\tilde{j}+1}{8(C+1)}]$ we dictate the death times by the geometry of the strands in the angular interval $[\pi\frac{\tilde{j}-1}{8(C+1)}+\pi,\pi\frac{\tilde{j}+2}{8(C+1)}+\pi]$. We do so by changing the $\tilde{\psi}$ coordinates of the strands in the interval, so that the death times for the interval $[\pi\frac{\tilde{j}}{8(C+1)},\pi\frac{\tilde{j}+1}{8(C+1)}]$ change, but the birth times in the interval $[\pi+\frac{\pi}{8(C+1)},2\pi-\frac{\pi}{8(C+1)}]$ remain the same, see Figure 14. This in particular ensures that we don’t introduce (extra) crossings in the vineyard that are not present in the closed braid we start with.
    Figure 14: Manipulating the $\tilde{\psi}$ changes the Death times (vertical section of Figure 1, not to scale).
    Between these intervals we interpolate between the different geometries of strands (again by changing the $\tilde{\psi}$ coordinates), which we can do because as mentioned the death times do not influence the topology of the closure of the braid appearing in the persistence diagram.
    Now let us consider a crossing of two strands in the interval $[\pi\frac{\tilde{j}}{8(C+1)},\pi\frac{\tilde{j}+1}{8(C+1)}]$ that are both not born first. As in intermezzo we denote with a little bit of abuse of notation by $b^{j,c}(t)$ and $b^{k,\tilde{c}}(t)$ (for $B$, or $b^{j,c}_{0}(t)$, $b^{k,\tilde{c}}_{0}(t)$,$b^{j,\tilde{c}}_{l}(t)$ and $b^{k,c}_{l}(t)$ respectively in the case of $\mathcal{M}$) both the birth times for these strands as well as the Morse critical points, where $c,\tilde{c}$ denote different connected components. We further stress that as in the intermezzo the $j$ and $k$ indices in $b^{j,c}(t)$ and $b^{k,\tilde{c}}(t)$ refer to order along the closed braid not the order of insertion. In the case where $l=1$ we further assume that neither $j$ nor $k$ corresponds to the birth with the very high birth value, that is the point where in non-extended persistence the second $1$-cycle that lives forever is born.
    We distinguish two different cases, one where $c=\tilde{c}$ and one where $c\neq\tilde{c}$. We start with the latter: We’ll focus on the (somewhat simpler) $B$ case, as the $\mathcal{M}$ case is virtually identical. Because the birth times correspond to Morse critical points on different connected components changing order does not influence the pairing between Morse critical points and birth and death in the persistence diagrams. This means that as long as the death times are distinct in $[\pi\frac{\tilde{j}}{8(C+1)},\pi\frac{\tilde{j}+1}{8(C+1)}]$ and the cycle born at $b^{j,c}(t)$ dies before the one born at $b^{k,\tilde{c}}(t)$ for one $t$ in this interval, then the cycle born at $b^{j,c}(t)$ dies before the one born at $b^{k,\tilde{c}}(t)$ for all $t$ in this interval and at continuous and distinct death times. Let us write $D^{J^{\prime},c}(t)$ and $D^{K^{\prime},\tilde{c}}(t)$ for the death times. This means that the persistence diagram (at level $t$ in the vineyard) contains the points $(b^{j,c}(t),D^{J^{\prime},c}(t))$ and $(b^{k,\tilde{c}}(t),D^{K^{\prime},\tilde{c}}(t))$, each of which locally describe a vine. This in turn implies that if (locally) $D^{J^{\prime},c}(t)^{\prime}>D^{K^{\prime},\tilde{c}}(t)$ then the vine $(b^{j,c}(t),D^{J^{\prime},c}(t))$ in the vineyard crosses under the vine $(b^{k,\tilde{c}}(t),D^{K^{\prime},\tilde{c}}(t))$ and the reversed order of death corresponds to an under crossing. This is achieved (as mentioned) by manipulating the $\tilde{\psi}$ coordinate. An easy way to achieve this if the component $c$ passes under $\tilde{c}$ is to push all strands of $c$ in and all strands of $\tilde{c}$ out and the reverse for an over pass. If one would like to repeat this discussion for $\mathcal{M}$ one only need to add a lower index 00 or $l$ respectively.
    We note that in the case where $l=1$, the birth with the very high birth value, that is the point where in non-extended persistence the second $1$-cycle that lives forever is born and was denoted by $m$ in the intermezzo, is separate (because of the high birth value) and therefore gives a disconnected circle in the closure of the vineyard.
    Now we consider the case there $c=\tilde{c}$. To simplify notation we will drop the index $c$ from the notation altogether. We again focus on the $B$ case, the case $\mathcal{M}$ is almost identical. If the death times satisfy (7) for each $t$ in $[\pi\frac{\tilde{j}}{8(C+1)},\pi\frac{\tilde{j}+1}{8(C+1)}]$, then there are points $(b^{j}(t),D^{J-1}(t))$ and $(b^{k}(t),D^{K}(t))$ in the persistence diagram. We can ensure that (7) holds by changing the $\tilde{\psi}$ coordinates as before, see Figure 14. Because the assumption (7) does not constrain the relative order of $b^{j}(t)$ and $b^{k}(t)$ nor the relative order of $D^{J-1}(t)$ and $D^{K}(t)$. This means that we can change the order of birth time (which occurs thanks to the crossing in the closed braid $B$) during the course of the interval $[\pi\frac{\tilde{j}}{8(C+1)},\pi\frac{\tilde{j}+1}{8(C+1)}]$ and we can fix the relative order of $D^{J-1}(t)$ and $D^{K}(t)$ as needed (which we can again do by manipulating $\tilde{\psi}$, see Figure 14). If $D^{J-1}(t)>D^{K}(t)$ in $[\pi\frac{\tilde{j}}{8(C+1)},\pi\frac{\tilde{j}+1}{8(C+1)}]$ then the $b_{j}$ vine crosses over the $b_{k}$ vine, while for the reversed order the $b_{j}$ vine crosses crosses under the $b_{k}$ vine. We refer to the appendix for an extensive example of this procedure in the case of the braid depicted in Figure 1. This means that regardless of if we have an over or under crossing we locally push the strands with the critical points $D^{I},D^{II},\dots,D^{J-2},D^{K+1},D^{K+2},\dots D^{N}$ in (towards the centre of the annulus) by a lot, and the strands with critical point $D^{J},\dots,D^{K-1}$ out (from the centre of the annulus) by a lot. This leaves the strands with $D^{J-1}$ and $D^{K}$ in the middle, and one may push the strand with $D^{J-1}$ a little out and the $D^{K}$ a little in if we want the $b^{j}$ vine segment to pass over the $b^{k}$ part of the vine, with the reverse pushing for the under crossing. This in particular shows that the braid $B$ can be faithfully reconstructed in the persistence diagram, although we also introduce some one extra loop per connected component of the link if $l\neq 1$ and two if $l=1$.
The only thing left to remark is that the conditions of Lemma 3.4 and Corollary 3.10 can always be satisfied thanks to Remark 3.13.
3.4 Vineyard monodromy
In this section we give the second main result of this paper, namely that every order monodromy can be found in a vineyard and in any dimension $l$ of $H_{l}$, with $l\leq d-2$. This result follows almost immediately from Theorem 1.2, however given that it answers one of the main open questions of Arya et al. [3] we still formulate it as a theorem.
See 1.1
Proof 3.15.
The result follows by applying Theorem 1.2 to the ($l$-offset) of the ouroborus knot, that is the closure of the braid with $k+1$ strands and $k$ over crossings, as depicted in Figure 13.
Appendix A Example
As an example of pushing outward and inward of the strands by manipulating the $\tilde{\psi}$ coordinates in Step 4 of the proof of Theorem 1.2 we discuss the example shown in Figure 1 in detail. We number the crossings $c_{1},\dots,c_{13}$ as indicated in Figure 15. The manipulation proceeds as below. We emphasize that all of the death values associated to points near a crossing need to be generic, that is no two identical death values. We will not repeat this for every crossing.
Figure 15: The crossings and the specific correspondence between the deaths or critical points and the strands indicated.
  • At $c_{1}$ yellow crosses under red, hence we push all red strands out and all yellow ones in. We will refer to this as yellow before red.
  • At $c_{2}$ the strand on which $b^{2}$ lies crosses under the strand with $b^{3}$, in accordance with (7) we push $D^{III}$ in and $D^{II}$ out so that $D^{I}$ lies in between (blue in the figure), what you do with the yellow strands doesn’t matter (as long as it is generic). The way that we push is also indicated in blue in the figure.
  • At $c_{3}$ red crosses under yellow, hence we push the yellow strands out and the red ones in, i.e. red before yellow.
  • At $c_{4}$ red crosses over yellow, hence push the red strands out and the yellow ones in, i.e. yellow before red.
  • Similarly to the crossing at $c_{2}$ (but at $c_{5}$, the strand with $b^{2}$ crosses under the strand $b^{3}$), at $c_{5}$ we push $D^{I}$ in and push $D^{II}$ out so that $D^{III}$ (green in the figure) lies in the middle (there is no condition on the yellow strands except genericity). Indicated in green in the figure.
  • At $c_{6}$ red goes before yellow.
  • The crossings $c_{7}$, $c_{8}$, and $c_{9}$ almost coincide in the figure, however, the red crosses in all cases over yellow, so that yellow needs to go before red [^3] Strictly speaking it is not necessary that yellow needs to go before red as long as you choose consistently for both crossings because at the red crossing the first birth changes from one strand to another, which leads to surgery so that the outer strand disconnects, which means that after a Reidemeister II move you are fine. However to fit with the text in the proof it is best to have yellow before red., because the first birth in red exchanges strand the coupling is automatic and leads to a disconnected component (i.e. surgery is performed).
  • At $c_{10}$ yellow goes before red.
  • At $c_{11}$ the order of the strands really doesn’t matter as long as all death values are distinct (generic), because this is another first birth interchange (on yellow this time).
  • The crossing at $c_{12}$ is again similar to the crossings at $c_{5}$ with the strand with $b^{2}$ crossing under $b^{3}$. We stress that because we label the critical points along the braid in order along the braid from the first birth and the first birth has exchanged strand the strand that contains $b^{2}$ is not the same strand as the one that contained $b^{2}$ at $c_{5}$ (where by the same we mean identification via shortest paths on the link). The strands that are associated to $D^{I}$, $D^{II}$, and $D^{III}$ change as a consequence as well, see Figure 15 (purple). This having been said, we follow the same procedure as at $c_{5}$: At $c_{12}$ push $D^{III}$ out and $D^{II}$ in, so that $D^{I}$ ends up in the middle (there is no condition on the yellow strands except genericity). Indicated in purple in the figure.
  • At $c_{13}$ red goes before yellow.